# URS: A Unified Neural Routing Solver for Cross-Problem Zero-Shot Generalization

### v1

@@ Abstract
Multi-task neural routing solvers have emerged as a promising paradigm for their ability to solve multiple vehicle routing problems (VRPs) using a single model. However, existing neural solvers typically rely on predefined problem constraints or require per-problem fine-tuning, which substantially limits their zero-shot generalization ability to unseen VRP variants. To address this critical bottleneck, we propose URS, a unified neural routing solver capable of zero-shot generalization across a wide range of unseen VRPs using a single model without any fine-tuning. The key component of URS is the unified data representation (UDR), which replaces problem enumeration with data unification, thereby broadening the problem coverage and reducing reliance on domain expertise. In addition, we propose a Mixed Bias Module (MBM) to efficiently learn the geometric and relational biases inherent in various problems. On top of the proposed UDR, we further develop a parameter generator that adaptively adjusts the decoder and bias weights of MBM to enhance zero-shot generalization. Moreover, we propose an LLM-driven constraint satisfaction mechanism, which translates raw problem descriptions into executable stepwise masking functions to ensure solution feasibility. Extensive experiments demonstrate that URS can consistently produce high-quality solutions for more than 100 distinct VRP variants without any fine-tuning, which includes more than 90 unseen variants. To the best of our knowledge, URS is the first neural solver capable of handling over 100 VRP variants with a single model.
@@ 1 Introduction
The Vehicle Routing Problem (VRP) is an essential class of combinatorial optimization problems (COPs) with extensive applications in logistics and supply chain management (Tiwari & Sharma, 2023; Sar & Ghadimi, 2023). Solving VRPs efficiently is challenging due to their NP-hard nature. While exact solvers can produce optimal solutions, they often become computationally prohibitive for real-world scenarios. In recent decades, classical heuristic solvers (Helsgaun, 2017; Vidal, 2022) have achieved impressive performance within acceptable timeframes. Nevertheless, these methods require considerable domain expertise to design specialized rules for each routing problem. Given the growing diversity of VRP variants in real-world applications, manually crafting tailored rules for every case has become impractical.
@@
In recent years, neural combinatorial optimization (NCO) methods have attracted substantial attention for their potential to reduce reliance on handcrafted rules while maintaining competitive solution quality (Bengio et al., 2021; Li et al., 2022a). This has led to the development of many high-performing neural routing solvers that can automatically learn implicit problem-specific patterns from data under different training paradigms such as supervised learning (SL) (Drakulic et al., 2023; Luo et al., 2023; 2025b), reinforcement learning (RL) (Bello et al., 2016; Kool et al., 2019; Zhou et al., 2024a), and emerging self-improved learning (SIL) (Luo et al., 2023; 2025a; Pirnay & Grimm, 2024). Although these solvers have demonstrated impressive performance on specific problems (Kwon et al., 2020; Kim et al., 2022; Zhou et al., 2025), they typically require architectural customization and per-problem retraining to accommodate the distinct constraints and features of different VRP variants. These limitations increase the overall training cost and hinder practical deployment on new problems.
@@
Table 1: Comparison between our URS and existing neural solvers with multi-task learning.
Neural Solver Training Paradigm Routing Problems (#\#) Zero-shot Generalization Remarks
MTPOMO (Liu et al., 2024a) RL 16 $\checkmark$ Constraint Combination
MVMoE (Zhou et al., 2024b) RL 16 $\checkmark$ Constraint Combination
RouteFinder (Berto et al., 2024) RL 48 $\checkmark$ Constraint Combination
CaDA (Li et al., 2025a) RL 16 $\checkmark$ Constraint Combination
MTL-KD (Zheng et al., 2025a) RL+KD 16 $\checkmark$ Constraint Combination
TSP-FT (Lin et al., 2024) RL 3 $\times$ Adapter-based Fine-tuning
MTL-MAB (Wang et al., 2025a) RL 3 $\times$ Adapter-based Fine-tuning
@@
GOAL (Drakulic et al., 2025) SL 9 $\times$ Adapter-based Fine-tuning
[…]
@@
To address the challenge of cross-problem generalization, growing attention has been directed toward multi-task learning capable of handling diverse routing problems. As summarized in Table 1, existing efforts largely fall into two categories: 1) constraint combination-based multi-task learning and 2) adapter-based fine-tuning frameworks. In the first approach, VRP variants are treated as different combinations of constraint attributes, and a unified model is trained across these combinations to enable knowledge sharing for problems with seen constraints (Liu et al., 2024a; Zhou et al., 2024b; Li et al., 2025a; Berto et al., 2024; Zheng et al., 2025a). The second approach builds a shared model backbone for all problems and incorporates specific input/output adapters for each problem, thereby reducing the retraining costs (Lin et al., 2024; Drakulic et al., 2025; Wang et al., 2025a).
@@
Despite the above advancements, existing methods still fall short in zero-shot generalization. The problem coverage of constraint combination-based approaches is inherently bounded by their manually specified constraint sets, while adapter-based approaches require fine-tuning and thus cannot perform zero-shot generalization. Moreover, both strategies fundamentally rely on the explicit enumeration of problems through a predefined set of problem tags. This practice is problematic because the constraint space of routing problems is open-ended and compositional, making any fixed tagging scheme fundamentally limited. More critically, creating and maintaining such a problem taxonomy requires considerable domain expertise, which is precisely what NCO always aims to avoid. A detailed review of related work on single-task and multi-task learning is provided in Appendix A.
In this paper, we propose a powerful Unified Neural Routing Solver (URS) to significantly improve the cross-problem zero-shot generalization ability for NCO methods. Unlike existing approaches, we introduce a unified data representation (UDR) that replaces problem enumeration with data unification, thereby broadening problem coverage and reducing reliance on domain expertise. In addition, we propose a Mixed Bias Module (MBM) to better learn the geometric and relational biases inherent in various problems. Based on UDR, we further develop a parameter generator that adaptively adjusts decoder and bias weights of MBM to improve zero-shot generalization. Furthermore, to handle complex and unseen routing problems, we introduce a constraint satisfaction mechanism that leverages a large language model (LLM) to translate raw problem constraint descriptions into executable node masking functions for each construction step, thereby ensuring our URS generates feasible solutions for various problems.
We evaluate URS on more than 100 VRP variants, including more than 90 unseen variants. The results demonstrate that URS not only obtains competitive performance against specialized neural solvers on seen variants but also exhibits strong zero-shot generalization capabilities on a wide range of unseen problems. To the best of our knowledge, URS is the first neural solver capable of efficiently solving over 100 distinct VRP variants using a single model without any retraining or fine-tuning.
@@ 2 Preliminaries
2.1 Vehicle Routing Problems
@@
Consider a VRP instance with an optional depot indexed by 0 and $n$ customers indexed by $\{1,2,\ldots,n\}$, such as in the Capacitated VRP (CVRP). The instance can be represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the node set $\mathcal{V}=\{v_{i}\}_{i=0}^{n}$ has a total size of $|\mathcal{V}|=1+n$ unless the variant has no depot (e.g., the Traveling Salesman Problem), and the edge set is $\mathcal{E}=\{e(v_{i},v_{j})\ |\ v_{i},v_{j}\in\mathcal{V},v_{i}\neq v_{j}\}$. The travel cost between nodes is given by a distance matrix $\bm{D}=\{d_{ij}|\ \forall\ i,j\in 0,\ldots,n\}$. In VRP, each node $v_{i}\in\mathcal{V}$ includes node coordinates $\{x_{i},y_{i}\}$ when available and problem-specific attributes (e.g., demands in CVRP). A feasible solution is a tour $\bm{\pi}=(\pi_{1},\pi_{2},\ldots,\pi_{m})$ (i.e., a finite sequence), which is a permutation of $m$ nodes. Let $\Omega$ denote the feasible sequence set that satisfies the constraints, given an objective function $f(\bm{\pi}|\mathcal{G})$, we aim to search for a sequence $\bm{\pi}^{*}$ with the maximum objective, i.e., $\bm{\pi}^{*}=\arg\max_{\bm{\pi}\in\Omega}f(\pi|\mathcal{G})$. In most routing problems, the objective can be represented as the negative value of the total distance of a feasible sequence $\bm{\pi}$.
@@
In this paper, we consider over 100 VRP variants which may simultaneously have one or more constraints from the following categories: (1) Capacity (C); (2) Open Route (O); (3) Backhaul (B); (4) Backhaul and Priority (BP); (5) Duration Limit (L); (6) Time Windows (TW); (7) Multi-Depot (MD); (8) Prize Collecting (PC); (9) Asymmetric (A); (10) Pickup and Delivery (PD). Detailed definitions of these constraints are provided in Appendix B.
2.2 Constructive Neural Routing Solver
@@
Most constructive NCO methods employ an encoder-decoder architecture (Luo et al., 2023; Kwon et al., 2020). Let learnable model parameters be denoted as $\bm{\theta}=\{\bm{\theta}_{enc},\bm{\theta}_{dec}\}$. Without loss of generality, we present the prevailing construction pipeline using AM (Kool et al., 2019). Given an instance $\mathcal{G}$, raw node features are first mapped by a linear projection to initial embeddings $H^{(0)}=\{\mathbf{h}_{i}^{(0)}\}_{i=0}^{n}$. They are then processed by an encoder $\bm{\theta}_{enc}$ with $L$ attention layers to produce refined node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$. At each decoding step $t$, the decoder $\bm{\theta}_{dec}$ sequentially selects a node to append to the current partial solution $\pi_{1:t-1}=(\pi_{1},\pi_{2},...,\pi_{t-1})$, where $\pi_{1},\pi_{t-1}\in\mathcal{V}$ are the first and last visited node, respectively. This sequential process continues until a complete solution is formed. To guarantee the feasibility of generated solutions, a masking function $\mathcal{M}_{t}$ sets the selection probabilities of the following nodes to -$\infty$ during the construction process: nodes that (1) are visited or (2) violate problem-specific constraints (e.g., capacity).
3 The Proposed Model: Unified Neural Routing Solver
@@
In this section, we propose a powerful Unified Neural Routing Solver (URS), which can significantly improve cross-problem zero-shot generalization for NCO methods. The framework is illustrated in Figure 1, and its key components are elaborated below.
Figure 1: The overview of our URS with a 44‑node APDTSP instance as an example. (a) The solving pipeline of URS; (b) The proposed mixed bias module that incorporates multiple geometric priors, enabling it to efficiently learn the geometric and relational biases inherent in various problems.
@@ 3.1 Unified Data Representation
Existing multi-task neural solvers (Liu et al., 2024a; Zhou et al., 2024b; Drakulic et al., 2025; Wang et al., 2025a) fundamentally rely on the explicit problem enumeration through a predefined set of problem tags, which is challenging because the constraint space of routing problems is open-ended and compositional, making any fixed tagging scheme fundamentally limited. More critically, creating and maintaining such a problem taxonomy requires considerable domain expertise, which is precisely what NCO always aims to avoid.
From a data perspective, despite the substantial diversity in constraints among routing problems, their instances share a common instance representation, e.g., all input instances under the current constraint combinations (Liu et al., 2024a; Zhou et al., 2024b) can be formulated as instances of either CVRP or CVRPTW. Leveraging this commonality, we propose a Unified Data Representation (UDR) to replace problem enumeration with data unification, thereby broadening problem coverage and reducing reliance on domain expertise. For an instance $\mathcal{G}$, we define the unified representation $\bm{U}=\{\mathbf{u}_{i}\}_{i=0}^{n}$, each $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$, where $\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}$ are a position identifier, unified attribute set, and node-type indicator, respectively.
Position Identifier ($\bm{\rho}_{i}$)
We define a unified positional identifier for each node as $\bm{\rho}_{i}=\{\eta_{i},x_{i},y_{i}\}\in[0,1]^{3}$. The $\eta_{i}\sim\mathrm{Uniform}(0,1)\in\mathbb{R}^{1}$ is an additional sampled scalar, which is used to address asymmetric problems, following  Drakulic et al. (2023). For symmetric cases, we simply set $\eta_{i}=0$.
@@
We define the unified attribute set as $\bm{\omega}_{i}=\{\delta_{i},\epsilon_{i},\mu_{i},\ e_{i},\ l_{i},\ s_{i}\}$, where they correspond to demand, prize, penalty, earliest arrival time, latest arrival time, and service time, respectively. The unification empowers the model to learn the relative importance of individual attributes in a shared embedding space. Furthermore, it supports effortless extension or ablation of attributes through simple zero-filling, which eliminates the need for changes to the model architecture.
@@
We provide an additional 5-dimensional multi-hot type tensor $\bm{\xi}_{i}\in\{0,1\}^{5}$ for each node, encoding: depot, pickup node, delivery node, node in sub-routes, node in open route. The nodes in sub-routes and open routes mean the solution $\pi$ can have sub-routes or be an open route. We introduce explicit structural roles via $\bm{\xi}_{i}$ that help the model generalize over diverse routing problems. Further ablation analysis concerning $\bm{\xi}_{i}$ can be found in Appendix K.1.
@@
Through the UDR, each problem instance activates only a specific subset of the unified feature space. Thus, for any instance $\mathcal{G}_{i}$ drawn from different problems, a corresponding multi-hot problem representation $\bm{\lambda}{i}$ is obtained by indicating active (non-zero) feature slots with a value of $1$. This representation captures only the presence of features, in contrast to their raw values, and it subsequently provides conditional guidance to the position bias weight (see Equation 6) and the adaptive decoder parameters $\bm{\theta}_{dec}(\bm{\lambda})$ (see Equation 8 and Equation 9), thereby helping the model in distinguishing problem variants and refining its node selection process.
Instead of relying on a predefined problem taxonomy with a discrete set of problem tags, our proposed UDR decouples instance representation from constraint definitions and operates in a unified space characterized by both continuous and discrete features. It allows a single neural solver to address a much larger, open-ended space of VRP variants, as new problems can be seamlessly incorporated into this unified representation. Problem-specific constraints are delegated to the model-agnostic masking function $\mathcal{M}_{t}$. For example, for all capacity-constrained problems, only the remaining load is explicitly provided as input, while diverse additional constraints (e.g., time windows, duration limit, and backhaul) are enforced implicitly by masking infeasible candidate nodes during solution construction, rather than being fully enumerated in the decoder’s input. This process is further automated by an LLM-driven constraint satisfaction mechanism that generates problem-specific mask generators, which convert constraint handling from manual engineering to automated programming (see Section 4 for more details). Together, our proposed UDR enables a single model to handle various explored and even unexplored VRP variants while significantly reducing reliance on domain expertise. Detailed descriptions of $\bm{U}$ and $\bm{\lambda}$ are provided in Appendix C and Appendix D, respectively.

### v2

@@ Abstract
Multi-task neural routing solvers have emerged as a promising paradigm for their ability to solve multiple vehicle routing problems (VRPs) using a single model. However, existing neural solvers typically rely on predefined problem constraints or require per-problem fine-tuning, which substantially limits their zero-shot generalization ability to unseen VRP variants. To address this critical bottleneck, we propose URS, a unified neural routing solver that achieves zero-shot generalization across a wide range of unseen VRPs with a single model. We propose a unified data representation (UDR) that replaces problem enumeration with data unification, thereby broadening the problem coverage and reducing reliance on domain expertise. In addition, we introduce a Mixed Bias Module (MBM) during encoding to improve node embeddings, which efficiently captures multiple priors inherent to various problems. On top of the UDR, we develop a problem-conditioned parameter generator to further improve zero-shot generalization. Extensive experiments show that URS consistently produces high-quality solutions for 110 VRP variants (including 99 unseen variants) while demonstrating impressive scalability to large-scale instances with up to 7000 nodes. To the best of our knowledge, URS is the first neural solver to handle over 100 VRP variants with a single model. Our code is available at https://github.com/CIAM-Group/URS.
Machine Learning, ICML
@@ 1 Introduction
The Vehicle Routing Problem (VRP) is an essential class of combinatorial optimization problems (COPs) with extensive applications in logistics and supply chain management (Tiwari and Sharma, 2023; Sar and Ghadimi, 2023). Solving VRPs efficiently is challenging due to their NP-hard nature. While exact solvers can produce optimal solutions, they often become computationally prohibitive for real-world scenarios. In recent decades, classical heuristic solvers (Helsgaun, 2017; Vidal, 2022) have achieved impressive performance within acceptable timeframes. Nevertheless, these methods require considerable domain expertise to design specialist rules for each routing problem. Given the growing diversity of VRP variants in real-world applications, manually crafting tailored rules for every case has become impractical.
@@
In recent years, neural combinatorial optimization (NCO) methods have attracted substantial attention for their potential to reduce reliance on handcrafted rules while maintaining competitive solution quality (Bengio et al., 2021; Ba et al., 2026). This has led to the development of many high-performing neural routing solvers that automatically learn implicit problem-specific rules from data under different training paradigms such as supervised learning (SL) (Drakulic et al., 2023; Luo et al., 2023, 2025b), reinforcement learning (RL) (Kool et al., 2019; Zhou et al., 2026), and self-improved learning (SIL) (Luo et al., 2025a; Pirnay and Grimm, 2024). Although these solvers have demonstrated impressive performance on specific problems (Huang et al., 2025; Zhou et al., 2025), they typically require architectural customization and per-problem retraining to accommodate the distinct constraints and features of different VRP variants, thereby increasing overall training costs and hindering practical deployment to new problems.
@@
Table 1: Comparison between our URS and existing neural solvers with multi-task learning. Note that ”#\#VRP Variants” and ”Generalizable Scale” refer to the total number of tested problems and the maximum scale reported in the original papers, respectively.
Multi-task Neural Routing Solver Training #\#VRP Variants Symmetric Asymmetric Pickup and Delivery Zero-shot Training Generalizable Remarks
Paradigm VRPs VRPs Problems Generalization Scale Scale
MTPOMO (Liu et al., 2024) RL 16 $\checkmark$ $\times$ $\times$ $\checkmark$ $100$ $1000$ Constraint Combination
MVMoE (Zhou et al., 2024) RL 16 $\checkmark$ $\times$ $\times$ $\checkmark$ $100$ $1000$ Constraint Combination
RouteFinder (Berto et al., 2025) RL 48 $\checkmark$ $\times$ $\times$ $\checkmark$ $100$ $1000$ Constraint Combination
CaDA (Li et al., 2025a) RL 16 $\checkmark$ $\times$ $\times$ $\checkmark$ $100$ $200$ Constraint Combination
MTL-KD (Zheng et al., 2025) RL+KD 16 $\checkmark$ $\times$ $\times$ $\checkmark$ $100$ $1000$ Constraint Combination
TSP-FT (Lin et al., 2024) RL 4 $\checkmark$ $\times$ $\times$ $\times$ $100$ $1000$ Adapter-based Fine-tuning
MTL-MAB (Wang et al., 2025a) RL 3 $\checkmark$ $\times$ $\times$ $\times$ $100$ $1000$ Adapter-based Fine-tuning
@@
GOAL (Drakulic et al., 2025) SL 10 $\checkmark$ $\checkmark$ $\times$ $\times$ $100$ $1000$ Adapter-based Fine-tuning
URS (Ours) RL 110 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 100 7000 Unified Data Representation
@@
To address the challenge of cross-problem generalization, growing attention has been directed toward multi-task learning capable of handling diverse routing problems. As summarized in Table 1, existing efforts largely fall into two categories: 1) constraint combination-based multi-task learning and 2) adapter-based fine-tuning. In the first approach, VRP variants are treated as different combinations of constraint attributes, and a unified model is trained across these combinations to enable knowledge sharing for problems with seen constraints (Liu et al., 2024; Zhou et al., 2024; Li et al., 2025a; Berto et al., 2025; Zheng et al., 2025). The second approach builds a shared model backbone for all problems and incorporates problem-specific input/output adapters, thereby reducing retraining costs (Lin et al., 2024; Drakulic et al., 2025; Wang et al., 2025a).
@@
Despite the above advancements, existing methods still fall short in cross-problem zero-shot generalization. The problem coverage of constraint combination-based approaches is inherently bounded by their manually specified constraint sets, whereas adapter-based approaches require additional fine-tuning and thus cannot perform zero-shot generalization. Furthermore, both strategies fundamentally rely on the explicit problem enumeration through a predefined set of problem tags. This practice is challenging because the constraint space of VRP variants is open-ended and compositional. More critically, creating and maintaining such a problem taxonomy requires considerable domain expertise, which NCO always aims to avoid. A detailed review of related work on single-task and multi-task learning is provided in Appendix A.
In this paper, we propose a powerful Unified Neural Routing Solver (URS) to improve the cross-problem zero-shot generalization ability for NCO methods. Our contributions can be summarized as follows: (1) We propose a unified data representation (UDR) that replaces existing problem enumeration with data unification, significantly broadening problem coverage while mitigating reliance on domain-specific expertise; (2) We introduce a Mixed Bias Module (MBM) during encoding to improve node embeddings, which efficiently captures multiple priors inherent to various problems while reducing architectural redundancy; (3) On top of the UDR, we develop a problem-conditioned parameter generator to further improve zero-shot generalization; (4) Extensive experiments on 110 VRP variants show that URS not only achieves competitive performance against specialist neural solvers on seen variants but also exhibits strong zero-shot generalization across 99 unseen variants. Notably, URS demonstrates impressive scalability to large-scale instances with up to 7000 nodes. To the best of our knowledge, URS is the first neural solver to efficiently solve over 100 VRP variants with a single model, without retraining or fine-tuning.
@@ 2 Preliminaries
Vehicle Routing Problems
@@
Consider a VRP instance with an optional depot indexed by 0 and $n$ customers indexed by $\{1,2,\ldots,n\}$, such as in the Capacitated VRP (CVRP). The instance can be represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the node set $\mathcal{V}=\{v_{i}\}_{i=0}^{n}$ has a total size of $|\mathcal{V}|=1+n$ unless the variant has no depot (e.g., the Traveling Salesman Problem), and the edge set is $\mathcal{E}=\{e(v_{i},v_{j})\ |\ v_{i},v_{j}\in\mathcal{V},v_{i}\neq v_{j}\}$. The travel cost between nodes is given by a distance matrix $\bm{D}=\{d_{ij}|\ \forall\ i,j\in 0,\ldots,n\}$. In VRP, each node $v_{i}\in\mathcal{V}$ includes node coordinates $\{x_{i},y_{i}\}$ when available and problem-specific attributes (e.g., demands in CVRP). Formally, a feasible solution is denoted by a node permutation $\bm{\pi}=(\pi_{1},\pi_{2},\ldots,\pi_{m})$ (Toth and Vigo, 2002) (i.e., a finite sequence) that represents a set of valid routes. Each route is traversed by a single vehicle originating and terminating at its own depot or starting node, while strictly satisfying all constraints. Let $\Omega$ denote the feasible sequence set, given an objective function $f(\bm{\pi}|\mathcal{G})$, we aim to search for a sequence $\bm{\pi}^{*}$ with the maximum objective, i.e., $\bm{\pi}^{*}=\arg\max_{\bm{\pi}\in\Omega}f(\pi|\mathcal{G})$. In most VRPs, $f(\bm{\pi}|\mathcal{G})$ can be defined as the negative value of the total distance of $\bm{\pi}$.
@@
In this paper, we consider 110 VRP variants which may simultaneously have one or more constraints from the following categories: (1) Capacity (C); (2) Open Route (O); (3) Backhaul (B); (4) Backhaul and Priority (BP); (5) Duration Limit (L); (6) Time Windows (TW); (7) Multi-Depot (MD); (8) Prize Collecting (PC); (9) Asymmetric (A); and (10) Pickup and Delivery (PD). Detailed definitions of these constraints are provided in Appendix B.
Figure 1: The pipeline of our URS for solving 110 VRP variants using a single model without any fine-tuning, illustrated on APDTSP.
Constructive Neural Routing Solver
@@
Most constructive NCO methods employ an encoder-decoder architecture for solution construction (Luo et al., 2023; Kwon et al., 2020). Let learnable model parameters be denoted as $\bm{\theta}=\{\bm{\theta}_{enc},\bm{\theta}_{dec}\}$. Without loss of generality, we present the prevailing autoregressive construction pipeline using AM (Kool et al., 2019). Given an instance $\mathcal{G}$, raw node features are first mapped by a linear projection to initial embeddings $H^{(0)}=\{\mathbf{h}_{i}^{(0)}\}_{i=0}^{n}$. They are then processed by an encoder $\bm{\theta}_{enc}$ with $L$ attention layers to produce refined node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$. At each decoding step $t$, the decoder $\bm{\theta}_{dec}$ sequentially selects a node to append to the current partial solution $\pi_{1:t-1}=(\pi_{1},\pi_{2},\ldots,\pi_{t-1})$, where $\pi_{1},\pi_{t-1}\in\mathcal{V}$ are the first and last visited node, respectively. This sequential process continues until a complete solution is formed. To guarantee the feasibility of generated solutions, a masking function $\mathcal{M}_{t}$ sets the selection probabilities of the following nodes to -$\infty$ during the construction process: nodes that (1) are visited or (2) violate problem-specific constraints (e.g., capacity).
3 Methodology
@@
In this section, as illustrated in Figure 1, we propose a powerful Unified Neural Routing Solver (URS), which significantly improves cross-problem zero-shot generalization for NCO methods. Its key components are elaborated below.
@@ 3.1 Unified Data Representation
Existing multi-task neural solvers (Zhou et al., 2024; Drakulic et al., 2025) fundamentally rely on explicit problem enumeration via a predefined set of problem tags, which is challenging because the constraint space of VRPs is open-ended and compositional. More critically, creating and maintaining such a problem taxonomy requires considerable domain expertise, which NCO always aims to avoid.
From a data perspective, despite the substantial diversity in constraints, their instances share a common structural representation. For existing constraint combination-based methods (Liu et al., 2024; Zhou et al., 2024), their evaluated problems can be formulated as CVRPTW instances. Leveraging this structural commonality, we propose a Unified Data Representation (UDR) $\bm{U}=\{\mathbf{u}_{i}\}_{i=0}^{n}$. By replacing explicit problem enumeration with data unification, our UDR broadens problem coverage and reduces reliance on domain expertise. Notably, the inputs to existing multi-task neural solvers based on either problem-specific adapters (Drakulic et al., 2025) or constraint combinations (Zhou et al., 2024) can be viewed as special subsets of our proposed UDR.
Instead of relying on a predefined problem taxonomy with a discrete set of problem tags, UDR decouples instance representation from constraint definitions and operates in a unified space characterized by both continuous and discrete features. It allows a single neural solver to address a much larger, open-ended space of VRP variants, as new problems can be seamlessly incorporated into this unified representation. Problem-specific constraints are delegated to the model-agnostic masking function $\mathcal{M}_{t}$. For example, for all CVRP variants, only the remaining load is explicitly provided as input, while diverse additional constraints (e.g., time windows) are enforced implicitly by masking infeasible candidate nodes during solution construction, rather than being fully enumerated in the decoder’s input. Thus, UDR enables a single model to handle various VRP variants while significantly reducing reliance on domain expertise. Specifically, for an instance $\mathcal{G}$, each $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$, where $\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}$ are a positional identifier, unified attribute set, and node-type indicator, respectively.
Positional Identifier ($\bm{\rho}_{i}$)
We define a unified positional identifier for each node as $\bm{\rho}_{i}=\{\eta_{i},x_{i},y_{i}\}\in[0,1]^{3}$. The $\eta_{i}\sim\mathrm{Uniform}(0,1)\in\mathbb{R}^{1}$ is a sampled scalar, which is used to address asymmetric problems, following  Drakulic et al. (2023). For symmetric cases, we simply set $\eta_{i}=0$.
@@
We define the unified attribute set as $\bm{\omega}_{i}=\{\delta_{i},\epsilon_{i},\mu_{i},\ e_{i},\ l_{i},\ s_{i}\}$, where they correspond to demand, prize, penalty, earliest arrival time, latest arrival time, and service time, respectively. The unification empowers the model to learn the relative importance of individual attributes in a shared embedding space. It also supports effortless attribute extension or ablation via simple zero-filling, eliminating the need to change the architecture.
@@
We provide a 5-dimensional binary vector $\bm{\xi}_{i}\in\{0,1\}^{5}$ for each node, encoding: depot, pickup node, delivery node, node in sub-routes, node in open route. The nodes in sub-routes and open routes mean the solution $\pi$ can have sub-routes or be an open route. We introduce explicit structural roles to enable the model to generalize over diverse variants. Further ablation analysis concerning $\bm{\xi}_{i}$ is provided in Appendix K.1.
@@
Through the UDR, each problem instance activates only a specific subset of the unified feature space. For any instance $\mathcal{G}_{i}$ drawn from different problems, a corresponding multi-hot problem representation $\bm{\lambda}_{i}$ is obtained by indicating active (non-zero) feature slots with a value of $1$. This representation captures only the presence of features, in contrast to their raw values, and it subsequently provides conditional guidance to the position bias weight (see Equation 6) and the adaptive decoder parameters $\bm{\theta}_{dec}(\bm{\lambda})$ (see Equation 9 and Equation 10), thereby helping the model in distinguishing problem variants and refining its node selection process.
A more detailed description of $\bm{U}$ and $\bm{\lambda}$ is provided in Appendix C and Appendix D, respectively.